# DEER: A Benchmark for Evaluating Deep Research Agents on Expert Report Generation

### v2

@@ Abstract
As large language models advance, deep research systems capable of generating expert-level reports through multi-step reasoning and evidence-based synthesis are emerging. However, evaluating such reports remains challenging. Existing benchmarks often lack systematic evaluation criteria, rely heavily on LLM-based judges that may miss issues requiring expert judgment, and verify only a limited subset of explicitly cited statements rather than report-wide factual reliability. To address these limitations, we introduce DEER, a benchmark for evaluating expert-level deep research reports. DEER comprises 50 report-writing tasks spanning 13 domains, along with an expert-grounded evaluation taxonomy with seven dimensions and 25 subdimensions, operationalized into 101 fine-grained rubric items. To improve evaluation consistency, DEER provides task-specific Expert Evaluation Guidance to support LLM-based judging. Complementing rubric-based assessment, we propose a document-level fact-checking architecture that verifies both cited and uncited claims and quantifies the quality and reliability of the supporting evidence. Experimental results show that DEER exhibits strong correlation with human expert judgments and yields interpretable diagnostics of system strengths and weaknesses.
DEER: A Benchmark for Evaluating Deep Research Agents
on Expert Report Generation
Janghoon Han1, Heegyu Kim1, Changho Lee1, Dahm Lee1, Min Hyung Park1, Hosung Song1, Stanley Jungkyu Choi1, Moontae Lee1,2, Honglak Lee1,3 1LG AI Research, 2University of Illinois Chicago, 3University of Michigan, Ann Arbor Correspondence: hanjh9439@gmail.com
@@ 1 Introduction
Driven by rapid advances in large language models (LLMs), automated deep research systems are emerging as a core technology in both academia and industry OpenAI (2025); Google (2025); Anthropic (2025); Yang et al. (2025); Li et al. (2025b); Huang et al. (2025); Li et al. (2025c). Unlike traditional information retrieval systems, these systems address complex research queries by decomposing them into multiple steps and dynamically seeking additional information based on intermediate results. Through this process, they integrate information from diverse sources and synthesize multiple perspectives to produce reliable, evidence-based research reports Xu and Peng (2025); Zhang et al. (2025); Java et al. (2025). As a result, deep research systems demonstrate strong performance even on challenging benchmark tasks.
@@
Early evaluations of deep research systems relied primarily on complex reasoning benchmarks Rein et al. (2023); Phan et al. (2025), which indirectly assessed information-gathering, hypothesis testing, and multi-step reasoning through task performance. Subsequently, deep web search QA benchmarks were introduced to more directly measure systems’ web browsing and information retrieval abilities—core capabilities of deep research—by evaluating multi-step search, information integration, and answer derivation Wei et al. (2025a); Krishna et al. (2025); Mialon et al. (2023); Chen et al. (2025); Gou et al. (2025). More recently, deep research report benchmarks have emerged to evaluate the quality of generated reports from multiple perspectives, moving beyond simple short-answer-based evaluation Li et al. (2025b); Du et al. (2025); Consult (2025); Coelho et al. (2025).
@@
Despite these advancements, existing methodologies for evaluating deep research systems continue to face important limitations when applied to expert-level reports. First, although prior work has proposed high-level dimensions for report evaluation, these frameworks do not comprehensively specify the diverse elements required for expert reporting. In particular, many approaches Du et al. (2025); Patel et al. (2025); Wang et al. (2025); Wan et al. (2025) rely on LLMs rather than human experts to define detailed subdimensions and concrete evaluation items, which can introduce inconsistencies and reduce evaluation reliability. Second, even when a fixed rubric is provided, evaluations that rely on LLM judges may fail to identify issues that require expert judgment. Third, current approaches to source verification are typically restricted to sections with explicit citation markers, leaving the factual reliability of claims across the full report insufficiently examined. Together, these limitations hinder comprehensive and reliable evaluation of deep research systems.
@@
To address these limitations, we propose DEER (the DEep research Expert Report benchmark), which provides comprehensive and reliable evaluation of deep research reports through 50 report-generation tasks spanning 13 domains. DEER analyzes a broad range of expert report assessment criteria and consolidates them via expert consensus into a Deep Research Report Evaluation Taxonomy comprising 7 dimensions and 25 subdimensions. Based on this taxonomy, DEER evaluates each report across two complementary components: (i) report-quality dimensions that require holistic, document-level judgment, and (ii) external-information dimensions that can be assessed at the claim level by fact verification. For report quality, we operationalize the taxonomy into a fixed set of 101 fine-grained rubric items and supplement these with report-specific Expert Evaluation Guidance authored by domain experts to improve the consistency and validity of LLM-based scoring. For external information, we introduce a fact-verification module that examines both cited and uncited claims across the report and produces quantitative measures of evidence quality. Finally, DEER integrates rubric-based LLM evaluations with the resulting metric-based signals to yield a unified, multi-dimensional assessment of each expert report. Figure 1 presents an overview of dimension-level performance for major deep research systems evaluated using DEER.
@@
  • We operationalize this taxonomy into a fixed set of 101 fine-grained rubric items and provide report-specific Expert Evaluation Guidance, improving the consistency and reliability of LLM-based evaluation.
  • We propose a report-level fact verification architecture that backtracks inter-claim dependencies to retrieve and verify relevant citations, enabling more complete evaluation of factual reliability across full reports.
@@ 2 Related Works
With the proliferation of high-performing LLMs, LLM-as-a-Judge approaches—using LLMs as evaluators—have been widely proposed and studied Zheng et al. (2023); Liu et al. (2023b); Chiang et al. (2024); Kim et al. (2024a, b). In this line of work, early evaluations for deep research primarily focused on how well models solve expert-level, high-difficulty questions Phan et al. (2025); Rein et al. (2023). Subsequently, to directly assess a core property of deep research—accessing and leveraging external information from the open web—benchmarks were proposed to measure models’ ability to browse/search the web, synthesize the required information, and construct answers grounded in that evidence Gou et al. (2025); Wei et al. (2025a); Huang et al. (2025). More recently, beyond simple short-answer evaluation, studies have sought to evaluate the ability to generate long-form reports that require expert-level analysis, reasoning, and interpretation by integrating information from multiple sources Consult (2025); Coelho et al. (2025); Du et al. (2025).
Despite this progress, existing report-evaluation methods are not based on expert-defined, fine-grained criteria. Some approaches rely on only a few coarse axes, leaving judgments largely to the evaluator LLM’s implicit standards Consult (2025); Coelho et al. (2025). Even when rubrics are subdivided, their specification and application can still depend substantially on the evaluator LLM Du et al. (2025), making it unclear whether the resulting scores correctly reflect report quality or align with expert assessment. In addition, source verification is often omitted Consult (2025); Coelho et al. (2025) or limited to a subset of cited claims Du et al. (2025), which may be insufficient to assess report-level factuality. To address these limitations, we propose an evaluation framework that (i) scores long-form research reports along multiple dimensions using expert-systematized, fine-grained criteria, and (ii) performs systematic source verification for claims throughout the report. A detailed comparison with prior work is provided in Appendix A.
@@ 3 Data Construction
To construct an evaluation dataset that reflects real-world usage, we analyze 5,842 in-house user queries collected from an internal deep research system and derive a topic distribution. For topic classification, we adopt the taxonomy of Wettig et al. (2025), consistent with prior work Du et al. (2025).
We use Humanity’s Last Exam (HLE) Phan et al. (2025) as the source of seed questions because it provides expert-written, high-difficulty, multidisciplinary questions that align well with the expert-level topics addressed in deep research reports. To match our target topic distribution, we map our topics to the 13 HLE subject domains and sample 50 seed questions accordingly.
@@ 4.1 Overview
To systematically evaluate deep research reports, we establish an evaluation framework grounded in two core aspects of deep research: external evidence acquisition and report-level synthesis Java et al. (2025); Xu and Peng (2025); Zhang et al. (2025). Based on these aspects, we construct a Deep Research Report Evaluation Taxonomy comprising seven major dimensions (§ 4.2), grouped into five report-quality dimensions and two external-information dimensions. The report-quality dimensions focus on synthesis and presentation through holistic, document-level judgment, whereas the external-information dimensions assess how external evidence is acquired and used via claim-level verification.
@@ 4.3 Expert-Guided Report Quality Assessment
Existing evaluation methods exhibit two key limitations when applied to long-form expert reports. First, they typically apply broad, high-level criteria, granting LLM judges substantial discretion over what aspects of a report to attend to. As a result, judges may (i) examine only parts of a report, (ii) consider only a subset of the many sub-aspects implied by each criterion, and (iii) focus on non-critical or superficial elements, leading to increased variance across judges and reduced evaluation consistency Li et al. (2025b); Consult (2025); Coelho et al. (2025). Second, even when the evaluation focus is clearly defined, assessing correctness and completeness often requires domain expertise that LLM judges may lack. Consequently, they may miss subtle but important issues, including non-obvious logical leaps, domain-specific misinterpretations, and fine-grained inaccuracies Du et al. (2025). To address these limitations, this study (1) makes evaluation criteria explicit through a fixed set of fine-grained rubric items and (2) provides task-specific Expert Evaluation Guidance that includes domain-specific context and reference points, enabling LLM judges to identify hard-to-detect errors and omissions.
Granular Rubric Design. Broad evaluation criteria can lead to inconsistent LLM-based scoring. Recent work addresses this issue by decomposing high-level criteria into more granular rubrics Lee et al. (2025); Wei et al. (2025b); Ruan et al. (2025). For deep research report evaluation, Du et al. (2025) follows a related approach by using an LLM to generate task-specific evaluation criteria. While this strategy narrows the evaluation focus, it raises two important concerns: whether the generated criteria are sufficiently comprehensive and capture what matters most in expert reports, and whether dynamically varying criteria enable interpretable and comparable scores that support systematic diagnosis of system strengths and weaknesses across tasks.
This study uses a fixed, expert-designed rubric to ensure reliable and interpretable evaluation. Building on the taxonomy in Table 9, a panel of experts operationalizes the 25 sub-dimensions into 46 evaluation criteria and translates each criterion into concrete, checkable rubric items. The items are organized into two aspects: coverage, whether required components are present and fully addressed wherever they occur in the report; and quality, the degree to which the targeted components are executed to a high standard. This process yields 101 scorable rubric items in total. [^1] The 46 criteria were validated through independent review by 10 experts following the procedure in § 4.2. The rubric is applied identically to all deep research reports across 50 tasks, enabling diagnosis of the system’s overall strengths and weaknesses with richer, more interpretable signals (see Appendix E for the rubric structure and examples).
@@
Expert Evaluation Guidance enumerates the required content elements and expert expectations for each task as concrete, verifiable statements that can be checked directly against the report. The guidance for each task is produced using the same expert drafting and cross-review procedure as the query reformulation process described in Section 3. Specifically, a domain expert reformulates each HLE item into a report-style prompt with explicit requirements (Section 3). The expert then derives the guidance by identifying the substantive content an adequate expert report should cover under the task prompt. Each required element is expressed as a concrete, verifiable statement that can be checked against the report. The resulting guidance is subsequently cross-reviewed by another expert from the same field to identify omissions, redundancies, or ambiguities, and revised as needed. Through this process, Expert Evaluation Guidance complements the fixed rubric by anchoring evaluation in task-specific expert knowledge while maintaining consistency and interpretability across tasks. Appendix B provides detailed guidance construction procedures and illustrative examples.
Model Req. Ful. Analyt. Sound. Struct. Coh. Format & Style Info. Int. Info. Suff. Ethics Mean
@@
Table 1: Baseline models’ expert report evaluation results (updated). In each column, the highest score is marked in bold and the second highest score is underlined.
@@ 4.4 Information Verification
To automate information verification in expert reports, we must overcome the limitations of existing systems that focus solely on explicit citations. In reality, many sentences implicitly inherit preceding evidence or construct new claims through contextual connections. Excluding these implicit claims from verification makes the reliability assessment of the entire document incomplete. Our Information Verification Module proposes an architecture that extends the verification scope to the entire document by tracking sentence-level logical flows.
Implicit Claim & Back-Tracking
This module analyzes contextual correlations to convert even uncited sentences into verifiable units. Specifically, when sentence $s_{i}$ implicitly cites the content of a previous sentence or paragraph, we use a Back-Tracking mechanism to trace the reference source. To this end, the LLM identifies the set of preceding sentences $R(s_{i})$ that $s_{i}$ depends on, and the valid citation set $\mathcal{V}(s_{i})$ used for actual verification is defined as follows:
$\mathcal{V}(s_{i})=\mathcal{C}(s_{i})\cup\bigcup_{k\in R(s_{i})}\mathcal{C}(s_{k})$ (1)
where $\mathcal{C}(s_{i})$ is the explicit citation of $s_{i}$. This structure allows verification by inheriting citations from referenced preceding sentences even if the sentence itself has no citation ($\mathcal{C}(s_{i})=\emptyset$). This is a key element that substantially expands the verification scope in policy and research reports containing many implicit claims.
Batch Processing for Claim Extraction
When processing long documents as a single input, LLMs suffer from the Lost-in-the-Middle Liu et al. (2023a) problem, which leads to the omission of important claims. To address this, we apply a Batch Extraction strategy that maintains the context of the entire document $D$ while parallel-processing only the target sentences divided into small batches $B_{j}$:
$\mathcal{A}=\bigcup_{j=1}^{M}E(D,B_{j}),\quad D=\bigcup_{j=1}^{M}B_{j}$ (2)
This increases the model’s focus to prevent missing important claims, and experimental results showed a significant improvement in extraction recall compared to the single processing ($E(D,D)$) method. Although the number of API calls increases, we confirmed that this is an essential strategy for securing high coverage in reports dealing with hundreds of claims.
Scalable Verification Architecture
The extracted claims are passed to the fact verification stage, where the biggest technical challenge is the trade-off between accuracy and cost. Verifying every claim individually yields high accuracy, but cost increases linearly with long documents. To address this, we introduce two scaling strategies. First, we apply Grouped Claim Verification, which bundles multiple claims citing the same source into a single input for verification. Experimental results showed that increasing Group Size from 1 to 20 raised F1 from 87.25 to 91.61 and reduced cost per 1K claims by 72%. This indicates that grouped input improves understanding the overall structure of the document and verifying claims, while reducing verification overhead. Second, greater API cost reduction is possible through the Context Retrieval strategy, which selectively provides only the context necessary for fact verification. This method achieved an API cost reduction of approximately 72%, but a 3.4% drop in F1 was observed, indicating a clear trade-off. This hierarchical pipeline (Batch Extraction $\rightarrow$ Grouped Verification $\rightarrow$ Context Retrieval) is designed to achieve both high-precision verification and cost efficiency even in long reports containing hundreds of claims.

### v3

@@ Abstract
Recent advances in large language models have enabled deep research systems that generate expert-level reports through multi-step reasoning and evidence-based synthesis. However, evaluating such reports remains challenging: report quality is multifaceted, making it difficult to determine what to assess and by what criteria; LLM-based judges may miss errors that require domain expertise to identify; and because deep research relies on retrieved evidence, report-wide claim verification is also necessary. To address these issues, we propose DEER, a benchmark for evaluating expert-level deep research reports. DEER systematizes evaluation criteria with an expert-developed taxonomy (7 dimensions, 25 subdimensions) operationalized as 101 fine-grained rubric items. We also provide task-specific Expert Evaluation Guidance to support LLM-based judging. Alongside rubric-based assessment, we propose a claim verification architecture that verifies both cited and uncited claims and quantifies evidence quality. Experiments show that while current deep research systems can produce structurally plausible reports that cite external evidence, there is room for improvement in fulfilling expert-level user requests and achieving logical completeness. Beyond simple performance comparisons, DEER makes system strengths and limitations interpretable and provides diagnostic signals for improvement. [^1] https://github.com/hanjanghoon/DEER.git
Machine Learning, ICML
@@ 1 Introduction
Driven by rapid advances in large language models (LLMs), automated deep research systems are emerging as a core technology in both academia and industry (OpenAI, 2025b; Google, 2025; Anthropic, 2025; Yang et al., 2025; Li et al., 2025b; Huang et al., 2025; Li et al., 2025c). Unlike conventional web search, these systems address complex research queries by decomposing them into multiple steps and dynamically seeking additional information based on intermediate results. Through this process, they integrate information from diverse sources and synthesize multiple perspectives to produce reliable, evidence-based research reports (Xu and Peng, 2025; Zhang et al., 2025; Java et al., 2025). As a result, deep research systems can achieve strong performance even on challenging benchmark tasks (Mialon et al., 2023; Phan et al., 2025).
@@
Early evaluations of deep research systems relied primarily on complex reasoning benchmarks (Rein et al., 2023; Mialon et al., 2023; Phan et al., 2025), which indirectly assessed information gathering, hypothesis testing, and multi-step reasoning through task performance. Subsequently, deep web search QA benchmarks were introduced to more directly measure systems’ web browsing and information retrieval abilities—core capabilities of deep research—by evaluating multi-step search, information integration, and answer derivation (Wei et al., 2025a; Krishna et al., 2025; Mialon et al., 2023; Chen et al., 2025; Gou et al., 2025). More recently, deep research report benchmarks have emerged to evaluate the quality of generated reports from multiple perspectives, moving beyond simple short-answer-based evaluation (Consult, 2025; Coelho et al., 2025; Du et al., 2025; Wan et al., 2025).
@@
Despite these advancements, existing methodologies for evaluating deep research systems continue to face important limitations when applied to expert-level reports. First, evaluation criteria are often underspecified, leaving it unclear what aspects of report quality should be assessed. Prior benchmarks typically evaluate reports using coarse, high-level dimensions, which do not provide sufficiently fine-grained criteria for precise assessment of report quality (Consult, 2025; Coelho et al., 2025). Moreover, even when fine-grained evaluation items are introduced, they are often generated or structured by LLMs, which can undermine consistency and reliability (Du et al., 2025; Wan et al., 2025; Wang et al., 2025). Second, even with well-specified rubric items, evaluations that rely on LLM judges may fail to identify issues that require domain expertise. Third, current approaches to source verification are typically restricted to claims with explicit citation markers, leaving factual reliability across the full report insufficiently examined. Together, these limitations hinder comprehensive and reliable evaluation of deep research systems.
@@
To address these limitations, we propose DEER (the DEep research Expert Report benchmark), which evaluates deep research reports through 50 report-generation tasks spanning 13 domains. DEER surveys established reporting norms and evaluation criteria across domains and synthesizes them, through an expert consensus process, into a Deep Research Report Evaluation Taxonomy comprising 7 dimensions and 25 subdimensions. Based on this taxonomy, DEER evaluates each report across two complementary components: (i) report-quality assessment, which requires holistic, document-level judgment, and (ii) external-information verification, which can be assessed at the claim level by checking against external sources. For report quality, we operationalize the taxonomy into a fixed set of 101 fine-grained rubric items and supplement them with task-specific evaluation guidance authored by domain experts to improve the consistency and validity of LLM-based scoring. For external information, we introduce an information-verification module that examines both cited and uncited claims across the report and produces quantitative measures of evidence quality. DEER then integrates rubric-based scores with these metric-based signals to yield a unified, multidimensional assessment of each expert report. Figure 1 presents an overview of dimension-level performance across deep research systems evaluated using DEER.
@@
  • We translate the taxonomy into a standardized set of 101 fine-grained rubric items and provide task-specific Expert Evaluation Guidance to support consistent and reliable LLM-based report scoring.
  • We propose a report-level information-verification architecture that backtracks inter-claim dependencies to retrieve citations and verify claims, enabling more complete evaluation of claim reliability across full reports.
@@ 2 Related Works
With the proliferation of high-performing LLMs, LLM-as-a-Judge approaches—using LLMs as evaluators—have been widely proposed and studied (Zheng et al., 2023; Liu et al., 2023b; Chiang et al., 2024; Kim et al., 2024a, b). In this line of work, early evaluations for deep research primarily focused on how well models solve expert-level, high-difficulty questions (Rein et al., 2023; Phan et al., 2025). Subsequently, to directly assess a core property of deep research—accessing and leveraging external information from the open web—benchmarks were proposed to measure models’ ability to browse/search the web, synthesize the required information, and construct answers grounded in that evidence (Gou et al., 2025; Wei et al., 2025a; Huang et al., 2025). More recently, beyond simple short-answer evaluation, studies have sought to evaluate the ability to generate long-form reports that require expert-level analysis, reasoning, and interpretation by integrating information from multiple sources (Consult, 2025; Coelho et al., 2025; Du et al., 2025; Wan et al., 2025; Wang et al., 2025; Yao et al., 2025; Sharma et al., 2025).
Despite this progress, existing report-evaluation methods are not based on expert-defined, fine-grained criteria. Some approaches rely on only a few coarse axes, leaving judgments largely to the evaluator LLM’s implicit standards (Consult, 2025; Coelho et al., 2025). Even when rubrics are subdivided, their specification and application can still depend substantially on the evaluator LLM (Du et al., 2025; Wan et al., 2025), making it unclear whether the resulting scores correctly reflect report quality or align with expert assessment. In addition, source verification is often omitted (Consult, 2025; Coelho et al., 2025) or limited to a subset of cited claims (Du et al., 2025; Wan et al., 2025), which may be insufficient to assess report-level factuality. To address these limitations, we propose an evaluation framework that (i) scores long-form research reports along multiple dimensions using expert-systematized, fine-grained criteria, and (ii) performs systematic source verification for claims throughout the report. A detailed comparison with prior work is provided in Appendix A.
@@ 3 Data Construction
To construct an evaluation dataset that reflects real-world usage, we analyze 5,842 in-house user queries collected from an internal deep research system and derive a topic distribution. For topic classification, we adopt the taxonomy of Wettig et al. (2025), consistent with prior work (Du et al., 2025).
We use Humanity’s Last Exam (HLE) (Phan et al., 2025) as the source of seed questions because it provides expert-written, high-difficulty, multidisciplinary questions that align well with the expert-level topics addressed in deep research reports. To match our target topic distribution, we map our topics to the 13 HLE subject domains and sample 50 seed questions accordingly.
@@ 4.1 Overview
To systematically evaluate deep research reports, we establish an evaluation framework grounded in two core aspects of deep research: external evidence acquisition and report-level synthesis (Java et al., 2025; Xu and Peng, 2025; Zhang et al., 2025). Based on these aspects, we construct a Deep Research Report Evaluation Taxonomy comprising seven major dimensions (§ 4.2), grouped into five report-quality dimensions and two external-information dimensions. The report-quality dimensions focus on synthesis and presentation through holistic, document-level judgment, whereas the external-information dimensions assess how external evidence is acquired and used via claim-level verification.
@@ 4.3 Expert-Guided Report Quality Assessment
Existing evaluation methods exhibit two key limitations when applied to long-form expert reports. First, they typically apply broad, high-level criteria, granting LLM judges substantial discretion over what aspects of a report to attend to. As a result, judges may (i) examine only parts of a report, (ii) consider only a subset of the many sub-aspects implied by each criterion, and (iii) focus on non-critical or superficial elements, leading to increased variance across judges and reduced evaluation consistency (Li et al., 2025b; Consult, 2025; Coelho et al., 2025). Second, even when the evaluation focus is clearly defined, assessing correctness and completeness often requires domain expertise that LLM judges may lack. Consequently, they may miss subtle but important issues, including non-obvious logical leaps, domain-specific misinterpretations, and fine-grained inaccuracies (Du et al., 2025). To address these limitations, this study (1) makes evaluation criteria explicit through a fixed set of fine-grained rubric items and (2) provides task-specific Expert Evaluation Guidance that includes domain-specific context and reference points, enabling LLM judges to identify hard-to-detect errors and omissions.
Granular Rubric Design. Broad evaluation criteria can lead to inconsistent LLM-based scoring. Recent work addresses this issue by decomposing high-level criteria into more granular rubrics (Lee et al., 2025; Wei et al., 2025b; Ruan et al., 2025). For deep research report evaluation, Du et al. (2025) follows a related approach by using an LLM to generate task-specific evaluation criteria. While this strategy narrows the evaluation focus, it raises two important concerns: whether the set of generated criteria is sufficiently comprehensive and captures what matters most in expert reports, and whether dynamically varying criteria enable interpretable and comparable scores that support systematic diagnosis of system strengths and weaknesses across tasks.
This study uses a fixed, expert-designed rubric to ensure reliable and interpretable evaluation. Building on the taxonomy in Table 9, a panel of experts operationalizes the 25 sub-dimensions into 46 evaluation criteria and translates each criterion into concrete, checkable rubric items. The items are organized into two aspects: coverage, whether required components are present and fully addressed wherever they occur in the report; and quality, the degree to which the targeted components are executed to a high standard. This process yields 101 scorable rubric items in total. [^2] The 46 criteria were validated through independent review by 10 experts following the procedure in § 4.2. The rubric is applied identically to all deep research reports across 50 tasks, enabling diagnosis of the system’s overall strengths and weaknesses with richer, more interpretable signals (see Appendix D for the rubric structure and examples).
@@
Expert Evaluation Guidance enumerates the required content elements and expert expectations for each task as concrete, verifiable statements that can be checked directly against the report. The guidance for each task is produced using the same expert drafting and cross-review procedure as the query reformulation process described in Section 3. Specifically, a domain expert reformulates each HLE item into a report-style prompt with explicit requirements (Section 3). The expert then derives the guidance by identifying the substantive content an adequate expert report should cover under the task prompt. Each required element is expressed as a concrete, verifiable statement that can be checked against the report. The resulting guidance is subsequently cross-reviewed by another expert from the same field to identify omissions, redundancies, or ambiguities, and revised as needed. Through this process, Expert Evaluation Guidance complements the fixed rubric by anchoring evaluation in task-specific expert knowledge while maintaining consistency and interpretability across tasks. Appendix B.4 provides detailed guidance construction procedures and illustrative examples.
Model Request. FulFill. Analyt. Sound. Struct. Cohere. Format & Style Inform Int. Inform. Suff. Ethics Mean
@@
Table 1: Evaluation results for expert reports generated by baseline methods. The best score in each column is shown in bold, and the second highest score is underlined.
@@ 4.4 Information Verification
To evaluate the external-information dimensions (Information Integrity and Information Sufficiency) in a reproducible way, we verify claims against evidence across the entire report and summarize the results as quantitative metrics, rather than relying on free-form LLM scoring. Specifically, our module (i) extracts atomic claims and links each verifiable claim to the sources it should be checked against (including uncited claims via implicit-citation recovery), and (ii) verifies whether the linked evidence supports each claim and aggregates the outcomes into Integrity/Sufficiency metrics (Appendix F.4).
Claim Types and Verification Scope.
Existing citation checkers typically verify only explicitly cited sentences, restricting verification to a narrow subset of report content. Following the human protocol (Appendix E), we classify extracted claims into six types (A–F) and focus external evidence verification on verifiable claim types (A–C) (Appendix F). This typing makes the verification scope explicit: (A) claims with inline citations; (B/C) uncited claims whose support is available elsewhere in the report; and (D–F) claims that are structural, non-verifiable, or lack identifiable support.
Implicit Claim Back-Tracking.
A key challenge is that expert reports often contain implicit claims whose supporting citations appear in earlier sentences (or even earlier sections), so the claim itself has no explicit marker. To verify such claims, we introduce a semantic Back-Tracking mechanism that recovers the citation set needed for verification. For a sentence $s_{i}$, the LLM identifies the set of preceding sentences $R(s_{i})$ that $s_{i}$ semantically depends on, and defines the valid citation set used for verification as:
$\mathcal{V}(s_{i})=\mathcal{C}(s_{i})\cup\bigcup_{k\in R(s_{i})}\mathcal{C}(s_{k}),$ (1)
where $\mathcal{C}(s_{i})$ denotes the explicit citations of $s_{i}$. This enables verification by inheriting citations from previously referenced sentences, even when $\mathcal{C}(s_{i})=\emptyset$, substantially expanding verification coverage beyond explicitly cited sentences.
Evidence-Grounded Verification and Metrics.
For each verifiable claim (Types A–C), we retrieve the evidence documents (URLs) cited in $\mathcal{V}(s_{i})$ and assess whether they support the claim under a strict support criterion (Appendix F). We then summarize claim-level outcomes into fine-grained metrics aligned with the Integrity/Sufficiency subdimensions (e.g., Claim Factuality, Citation Support, Evidence Coverage, and Reference Reliability/Diversity) and aggregate them into the two dimension-level scores used in our final evaluation (Appendix F.4). Implementation details for long-document processing and efficiency-oriented engineering (e.g., batching and grouping) are provided in Appendix F.4.